# DiTSE: High-Fidelity Generative Speech Enhancement via Latent Diffusion Transformers

### v1

@@ III-C Three-axis Evaluations
We compare the proposed methods to state-of-the-art baselines using objective and subjective evaluations. We consider three evaluations axes: (1) content accuracy measured by WER; (2) acoustic quality measured by quality mean-opinion-score (MOS); (3) speaker preservation measured by similarity MOS. We examine the methods on five evaluation sets unseen during training: DAPS Real, AQECC and DEMO, as utilized in prior work [5], each consisting of 100 samples representing diverse real-world scenarios; VoiceBank-DEMAND (VBD) [42] for denoising benchmark; and EARS [43] for challenging denoising and de-reverberation scenarios. For VBD and EARS, we sample twenty distinct speakers in each category with two samples per speaker, ensuring a variety of voices with balanced genders. We compare with the following methods: HiFi-GAN-2 [37], a GAN-based method for high-quality enhancement at ; transcription-free Miipher [40], a generative method which re-synthesizes clean speech from enhanced SSL features at ; Genhancer [5], a discrete token generation method at that is closest to our method in conditioning designs; SGMSE+ [7] at and StoRM [11] at as diffusion baselines. For generative methods, we evaluate the samples of the inference run with the median WER from five random runs. We use audio samples at the maximum sample rates offered by the models and the evaluations sets. We additionally include and down-sampled versions of our results for fair comparisons to the baselines at their sample rates.
@@
We conducted MOS listening tests via Prolific333https://www.prolific.co/. to evaluate audio quality and speaker similarity. For quality, listeners rated audio samples from 1 (Bad) to 5 (Excellent) in consideration of cleanness and naturalness, using a fixed clean DAPS sample as the reference of studio audio quality. For speaker similarity, listeners rated the preservation of the speaker’s voice in an audio sample with respect to the corresponding input audio, using the same scale. Each method received at least 400 ratings for each evaluation set in each test. More details on the listening test setup can be found on our demo page.
@@
Table II reports the results for the three evaluation axes. DiTSE stands out as the only model capable of reducing content error compared to the noisy input on DAPS and VBD, significantly outperforming all the baselines. EARS is particularly challenging as its spoken content is barely intelligible due to severe corruption by noise and reverberation. DiTSE is among the best of the generative methods on EARS, while HiFi-GAN-2 achieves slightly higher content accuracy but lower enhanced quality by manipulating the original waveform. Striking the balance between enabling content restoration and refraining models from hallucinations still remains an open problem under extreme SNR scenarios.
@@
Our methods significantly improve acoustic quality over all the baselines. To the best of our knowledge, DiTSE-Base is the first method that achieves a Quality MOS of 4.34 indistinguishable from the real studio-quality clean audio on DAPS. It also achieves an MOS comparable with Genhancer on AQECC, but under-performs on DEMO and VBD. We pinpoint that the VAE could limit audio quality as the reconstruction of clean audio (VAE Reconst.) displays notable reduction in Quality MOS from the ground-truth clean audio on VBD. Incorporating the post-hoc quantization (DiTSE-Base + Post) improves audio fidelity, achieving state-of-the-art audio quality across all the evaluation sets, while the content remains largely unchanged aside from a few subtle pronunciation variations at difficult spots (<0.5% WER increase).
@@
Lastly, Speaker MOS shows significant advantages of DiTSE in preserving speaker and paralinguistic features of the original speech over baselines. The robust conditioning enables the model to effectively attend to speaker traits in complex acoustic scenarios. Furthermore, generation in the continuous space facilitates the preservation of spoken nuances such as breath and coarticulation, resulting in more natural sounding compared to discrete token generation, while quantization after generation could further enhance fidelity of the voice.
IV Conclusions
@@
We introduce DiTSE, a speech enhancement approach built upon a latent diffusion transformer model. DiTSE effectively preserves spoken content and speaker characteristics by incorporating robust conditioning features. We additionally leverage post-hoc quantization to mitigate audio fidelity issues of diffusion-based approaches. Comprehensive evaluations demonstrate that our proposed approach efficiently achieves state-of-the-art audio quality comparable to real studio quality in 32 generation steps, while significantly improving content accuracy and speaker consistency over previous methods. Diffusion distillation techniques [44] could further reduce the computational costs of our approaches in the future.

### v2

@@ 3.3 Three-axis Evaluations
We compare the proposed methods to state-of-the-art baselines using objective and subjective evaluations. We consider three evaluation axes: (1) content accuracy measured by WER; (2) acoustic quality measured by quality mean-opinion-score (MOS); (3) speaker preservation measured by similarity MOS. We examine the methods on five evaluation sets unseen during training: DAPS Real, AQECC, and DEMO, as utilized in prior work [4], each consisting of 100 samples representing diverse real-world scenarios; VoiceBank-DEMAND (VBD) [36] for denoising benchmark; and EARS [37] for challenging denoising and de-reverberation scenarios. For VBD and EARS, we sample twenty distinct speakers in each category with two samples per speaker, ensuring a variety of voices with balanced genders. We compare with the following methods: HiFi-GAN-2 [33]; transcription-free Miipher [34]; Genhancer [4], a discrete token generation method that is closest to our method in conditioning designs; SGMSE+ [5] and StoRM [9] as diffusion baselines. For consistency, all baseline models are trained using the same dataset as our model (section 3.1).
@@
We conducted MOS listening tests via Prolific333https://www.prolific.co/. to evaluate audio quality and speaker similarity. For quality, listeners rated samples from 1 (Bad) to 5 (Excellent) for cleanness and naturalness, using a fixed clean DAPS sample as the reference of studio audio quality. For speaker similarity, listeners rated the preservation of the speaker’s voice in a sample with respect to the corresponding input, using the same scale. Each method received at least 400 ratings for each evaluation set.
We use the highest available sampling rates for each model and dataset. We also include and downsampled versions of our results for fair comparison to confirm that performance differences are not driven by sampling rates. Detailed results per sampling rate are available on our demo page.
@@
Table 2 reports the results for the three evaluation axes. DiTSE stands out as the only model capable of reducing content error compared to the noisy input on DAPS and VBD, significantly outperforming all the baselines. EARS is a challenging dataset with severe noise and reverberation and barely intelligible speech. It also includes complex out-of-domain sound, such as fillers, emotional interjections and laughter. DiTSE ranks among the top generative methods on EARS, while the regression-based HiFi-GAN-2, as it modifies the original waveform, yields slightly higher content accuracy but noticeably lower perceptual quality. Striking the right balance between enabling content restoration and refraining models from hallucinations still remains an open problem under extreme SNR scenarios.
@@
Our methods significantly improve acoustic quality over all the baselines. To the best of our knowledge, DiTSE-Base is the first method that achieves a Quality MOS of 4.34 that is indistinguishable from the real studio-quality audio of DAPS. It achieves an MOS comparable with Genhancer on AQECC, but under-performs on DEMO and VBD. We pinpoint that the VAE could limit audio quality as the reconstruction of clean audio (VAE Reconst.) displays notable reduction in Quality MOS on VBD. Incorporating the post-hoc quantization (DiTSE + Post) improves audio fidelity by regularizing harmonics of the generated speech and eliminating generation artifacts, achieving state-of-the-art audio quality across all the evaluation sets. The content remains largely unchanged aside from a few subtle pronunciation variations at difficult spots.
@@
Lastly, Speaker MOS shows significant advantages of DiTSE in preserving speaker and paralinguistic features over baselines. The robust conditioning enables the model to effectively attend to speaker attributes in complex acoustic scenarios. Furthermore, generation in the continuous space facilitates the preservation of spoken nuances such as breath and coarticulation, resulting in more natural speech compared to generative methods using discrete tokens, while quantization after continuous-space generation enhances voice fidelity.
4 Conclusions
@@
We introduce DiTSE, a speech enhancement approach built upon a latent diffusion transformer model. DiTSE effectively preserves spoken content and speaker characteristics by incorporating robust conditioning features. We additionally leverage post-hoc quantization to mitigate audio fidelity issues of diffusion-based approaches. Comprehensive evaluations demonstrate that our proposed approach efficiently achieves state-of-the-art audio quality, for the first time comparable to real studio quality, in 32 generation steps, while significantly improving content accuracy and speaker consistency over previous methods. Diffusion distillation techniques [38] could help reduce the computational cost of DiTSE in the future.